# Dealing with Logs and Zeros in Regression Models

### v2

@@ 2.5 Partialling-out fixed-effects
More generally, we view the simplifying assumption $E[M_{0}X_{1}\mid X_{0}]=0$ as natural whenever the outcome $Y_{i}$ does not call for an interactive fixed-effects specification. Should $X_{1}$ also exhibit genuine latent factors, one may replace the standard within-operator $M_{0}$ with a factor-purged projector, following Bai (2009) or Phillips and Williams (2024). [^13] For instance, the best rank-1 approximation of the double-demeaned matrix can be obtained via singular-value decomposition or an alternating least-squares routine. Although we do not pursue this extension here, our algorithm is readily adapted to such settings. For clarity and consistency, all simulations and applications presented below remove all fixed-effect dimensions, thereby abstracting from the possibility of remaining interaction terms.
@@ 6 Conclusion
Appendix A Proof of Theorem 1
Proving Theorem 1 requires multiple steps.
We first prove consistency and normality for the biased case, i.e. $\forall\delta<\infty$. Then we prove consistency and normality for the exact case, i.e. $\delta=\infty$. Theorem 1 is just a sequential application of these results.
@@ Appendix A Proof of Theorem 1
where $\tilde{Y}_{i}(\beta)=\log(Y_{i}+\delta\exp(X_{i}^{\prime}\beta))-c(\beta,\delta)$ is the transformed dependent variable. To further simplify exposition, we focus on $\delta=1$ in our calculations. The mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$ which characterizes the parameter is hence defined $\forall\phi\in\mathbb{R}^{K}$ as
@@
Our proof follows Dominitz and Sherman (2005), hereafter denoted DS, who develop a convergence theory for iterative estimators. Following DS, the convergence of iOLS requires that $M(\cdot)$ and $\hat{M}_{n}(\cdot)$ be contraction mappings, asymptotically. [^24] The reader is referred to DS for a formal definition of an asymptotic contraction mapping.
In order to show the convergence result $n^{1/2}|\hat{\beta}_{t(n)}-\beta|=O_{p}(1)$ as $n\to\infty$ by applying Theorem 1 in DS, we need to show that the following conditions hold:
@@
We verify conditions (i)-(iv) in turn. For (i), we adapt Dominitz and Sherman (2005) to show that $\{\hat{M}_{n}(\cdot)\}$ is an asymptotic contraction on $(B_{0},E_{K})$. In addition, we also show that, for any finite $n>0$ and any starting value $\phi$, $\hat{M}_{n}$ is a (finite-sample) global contraction. We report both results because the final debiasing step enjoys only a local contraction property, for which the asymptotic notion in Dominitz and Sherman (2005) is needed.
@@
Let us denote $X_{1i}=1$, for all $i$ as the constant. By prewhitening, we have $\sum_{j=1}^{n}X_{1j}=n$ and $\sum_{j=1}^{n}X_{kj}=0$ for $k>1$.
for $k>1$ and $\frac{\partial\hat{c}(\phi)}{\partial\phi_{1}}=0$. This expression simplifies when evaluated at $\phi=\beta$, as shown by
@@
for $k>1$ because $\hat{\phi}^{1}(\beta)=\log(n^{-1}\sum_{i=1}Y_{i}\exp(-X_{i}^{r^{\prime}}\beta^{r}))=\beta_{1}+\log(n^{-1}\sum_{i=1}U_{i})$, where $\log(n^{-1}\sum_{i=1}U_{i})=O_{p}(1)$ by iid assumption and $E[U_{i}]=1$, and $n^{-1}\sum_{i=1}^{n}X_{ki}=0$ by prewhitening. Thus, we have $\frac{\partial\hat{\phi}_{1}(\beta)}{\partial\phi_{k}}=0$.
Therefore, each element $(k,l)$ of $\hat{V}_{n}(\beta)$ writes
$\left[\hat{V}_{n}(\beta)\right]_{k,l}$ $=n^{-1}\sum_{i=1}^{n}\frac{X_{ki}X_{li}}{1+U_{i}}-n^{-1}\sum_{i=1}^{n}X_{ki}n^{-1}\sum_{j=1}^{n}\frac{X_{lj}}{1+U_{j}},$ (33)
for $l>1$ and
$\left[\hat{V}_{n}(\beta)\right]_{k,l}$ $=n^{-1}\sum_{i=1}^{n}\frac{X_{ki}}{1+U_{i}},$ (34)
for $l=1$. Remark that for $k=1,\forall l>1$ we have $\left[V_{n}(\beta)\right]_{1,l}=0$, and for $k=1,l=1$, we have $\left[\hat{V}_{n}(\beta)\right]_{1,1}=n^{-1}\sum_{i=1}^{n}\frac{1}{1+U_{i}}<1$. Therefore, the eigenvalue associated with the constant term is strictly below $1$, and proving the convergence amounts to showing that the largest eigenvalue of the $(K-1)\times(K-1)$ lower right-hand submatrix of $\hat{V}_{n}(\beta)$ is strictly less than unity. All elements $(k,l)$ for $k,l>1$ of this matrix writes
$\left[\hat{V}_{n}(\beta)\right]_{k,l}$ $=n^{-1}\sum_{i=1}^{n}\frac{X_{ki}X_{li}}{1+U_{i}}.$ (35)
@@
where $W$ is a diagonal matrix with elements $(i,i)$ acting as weights given by $\frac{1}{1+U_{i}}\in(0,1]$. Note that those weights become $\frac{\delta}{\delta+U_{i}}\in[0,1)$ for $\delta\neq 1$, thus the modulus of the contraction mapping will approach 1 as $\delta$ grows, meaning that smaller $\delta$ will lead to faster numerical convergence of the iterative algorithm.
We can thus write $W=W^{1/2}W^{1/2}$, and rewrite the submatrix of interest as the quadratic form
$\left[\hat{V}_{n}(\beta)\right]_{k,l>1}$ $=n^{-1}X^{\prime}W^{1/2}W^{1/2}X.$ (37)
Consequently, this matrix is nonnegative definite and so must have all nonnegative eigenvalues. We can alternatively write the weight matrix $W=I_{n}-D$, where $D$ is also a diagonal matrix with elements $\frac{U_{i}}{1+U_{i}}\in[0,1)$, or more generally $\frac{U_{i}}{\delta+U_{i}}\in[0,1)$. Therefore, we have the alternative expression
$\left[\hat{V}_{n}(\beta)\right]_{k,l>1}$ $=n^{-1}X^{\prime}(I_{n}-D)X=I_{K-1}-n^{-1}X^{\prime}D^{1/2}D^{1/2}X,$ (38)
where the second term is also a quadratic form. It follows that as $n\to\infty$, the maximum eigenvalue is equal to
$\max_{|a|=1}a^{\prime}\left[\hat{V}_{n}(\beta)\right]_{k,l>1}a=\max_{|a|=1}1-a^{\prime}X^{\prime}D^{1/2}D^{1/2}Xa.$ (39)
Assuming the data distribution is non-degenerate, $a^{\prime}X^{\prime}D^{1/2}D^{1/2}Xa$ is positive and bounded away from zero for all unit vectors $a\in\mathcal{R}^{K-1}$. Thus, as $n\to\infty$, the maximum eigenvalue of $\hat{V}_{n}(\beta)$ is strictly less than unity. This proves the result.
@@
where $W$ denotes the diagonal matrix with elements $\frac{\exp(X_{i}^{\prime}\phi)}{Y_{i}+\exp(X_{i}^{\prime}\phi)}\in(0,1]$. Therefore, the same logic applies and the algorithm is a contraction mapping for any $n$ and any $\phi$.
@@
Following the same reasoning as in the previous theorem, a sufficient condition for convergence is that $\frac{U_{i}}{1+\rho}$ is between 0 and $1$ for all $i$. Therefore, the choice of $\rho$ will not only affect the speed of convergence, but also whether the estimator converges at all.
@@
has element $1-\frac{U_{i}}{1+\delta}$, and $\hat{\Sigma}_{0}$ is an estimator of the covariance of $X_{i}^{\prime}U_{i}$ across observations.
Appendix B Proof of Theorem 2: 2SLS

### v3

@@ 2.5.2 All fixed-effects dimensions are short
Should $X_{1}$ also exhibit genuine latent factors, one may replace the standard within-operator $M_{0}$ with a factor-purged projector, following Bai (2009) or Phillips and Williams (2024), to partial out multiple fixed-effect dimensions. [^14] For two-way fixed-effects, the best rank-1 approximation of the double-demeaned matrix can quickly be obtained via singular-value decomposition. For larger dimensions, an alternating least-squares routine must be used. Although we do not pursue this extension here, our algorithm is readily adapted to such settings.
More generally, we view the simplifying assumption $E[M_{0}X_{1}\mid X_{0}]=0$ as natural whenever the outcome $Y_{i}$ does not call for an interactive fixed-effects specification. For simplicity, all simulations and applications presented in this paper remove all fixed-effect dimensions using this algorithm, thereby abstracting from the possibility of remaining interaction terms.
2.5.3 A mix of large and short fixed-effects dimensions
In the presence of a mix of large and short fixed-effects dimensions, we can combine both approaches. We can estimate the large dimensions of the fixed effects (e.g., time effects) using Algorithm 3 while applying Algorithm 4 only to the small dimensions (e.g., unit effects). [^15] With multiple small dimensions, again a factor-purged projector will be needed. In that setup, the transformed error no longer contains cross-terms between units and periods and therefore retains the same exogeneity properties as in the one-way case. [^16] For instance, in a panel with $T=2$ and $n=100$, one should estimate the time dummies and remove the unit dummies, which would otherwise be poorly estimated given only two time periods. This leads to the following procedure.
Let us partition the regressor matrix as $X=[X_{0},X_{1},X_{2}]$, where $X_{0}$ comprises the fixed-effect dummies for the large dimensions, $X_{1}$ contains the covariates (including the intercept), and $X_{2}$ contains the the fixed-effect dummies for the small dimensions to be partialled-out from $\Lambda$. We use the linear projections $P_{0}$ (onto $span(X_{0})$) and $\widetilde{P}_{2}$ (onto $span(M_{0}X_{2})$) and their orthogonal complements $M_{0}$ and $\widetilde{M}_{2}$. We define the residualized aggregate fixed-effect term by $\tilde{\Lambda}=\widetilde{M}_{2}X_{0}^{\prime}\beta_{0}$.
Algorithm 5 (iOLS-FD-HDFE)
[…]
  1. 0.
    Partial-out $X_{1}$ into $\acute{\acute{X_{1}}}=M_{0}\widetilde{M}_{2}X_{1}$;
  2. 1.
    Initialize $t=0$ and let $\hat{\beta}_{0}$ and $\hat{\tilde{\Lambda}}_{0}$ be initial estimates;
  3. 2.
    Transform $Y$ into $\tilde{Y}=\log(Y+\delta\exp(\widetilde{M}_{2}X_{1}\hat{\beta}_{t}+\hat{\tilde{\Lambda}}_{t}))-c(\delta,\hat{\beta}_{t},\hat{\tilde{\Lambda}}_{t})$;
  4. 3.
[…]
  5. 4.
    Run OLS: $\hat{\beta}_{t+1}=[\acute{\acute{X_{1}}}^{\prime}\acute{\acute{X_{1}}}]^{-1}\acute{\acute{X_{1}}}^{\prime}\acute{\tilde{Y}}$;
  6. 5.
    Recover: $\hat{\tilde{\Lambda}}_{t+1}=P_{0}\widetilde{M}_{2}\left(\tilde{Y}-X_{1}\hat{\beta}_{t+1}\right)$, and update $t$ to $t+1$;
  7. 6.
    Iterate steps (2.) to (5.) until $\hat{\beta}_{t}$ converges. Call the result $\hat{\beta}_{\delta}$.
  8. 7.
    Re-initialize $t=0$, update $\delta(i)$ to $\delta(i+1)$ and iterate steps (2.) to (7.) until $\hat{\beta}_{\delta(i)}$ converges. Call the solution $\hat{\beta}_{\delta}$. Proceed to Phase 2 to eliminate the remaining approximation bias $O(\delta^{-2})$ as in Algorithm 1.
The resulting estimator is based on the transformed model.
$M_{0}\widetilde{M}_{2}\log\left(\frac{Y_{i}+\delta\exp(\widetilde{M}_{2}X_{1i}\beta+\widetilde{M}_{2}\Lambda)}{\exp(c_{i}(\beta,\delta))}\right)=M_{0}\widetilde{M}_{2}X_{1i}\beta+M_{0}\widetilde{M}_{2}\overline{\upsilon}_{i},$ (20)
and the observation that $\widetilde{M}_{2}\Lambda$ can be recovered using the same approach as for iOLS-HDFE noting that $P_{0}\widetilde{M}_{2}=P_{0}$ by construction, so again we need $P_{0}\overline{\upsilon}_{i}$ to vanish asymptotically with the sample size.
To ensure global convergence of the doubly residualized map, the overlap requirement used for iOLS-FD must be strengthened to Assumption 8. For iOLS-FD, it suffices that there is positive overlap on $C$, i.e. a uniform slack away from $\tau_{i}=1$ along directions spanned by $M_{0}X_{1}$. By contrast, with the aggregate fixed-effects term the Jacobian contains the resolvent $(I-P_{0}T)^{-1}$ which can amplify weight concentration within FE groups.
Assumption 8 (Sufficient slack)
Let $\acute{\acute{X}}_{1}:=M_{0}\widetilde{M}_{2}X_{1}$ and $C:=\mathrm{col}(\acute{\acute{X}}_{1})\subset\mathbb{R}^{n}$. Define the diagonal weight matrix $T(\beta):=\mathrm{diag}\big(\tau_{1}(\beta),\ldots,\tau_{n}(\beta)\big)$ by
$\tau_{i}(\beta)\ :=\ \delta\,\exp\!\big((M_{0}\widetilde{M}_{2}X_{1}\beta)_{i}\big)\left(Y_{i}+\delta\,\exp\!\big((M_{0}\widetilde{M}_{2}X_{1}\beta)_{i}\big)\right)^{-1}\in(0,1]$, so $\tau_{i}(\beta)=1$ iff $Y_{i}=0$, and $\tau_{i}(\beta)\in(0,1)$ otherwise. There exist constants $\eta_{C}\in(0,1)$ and $\kappa_{\mathrm{FE}}\in[0,1)$ such that, uniformly over $\beta\in B_{0}$, $\eta_{C}\;>\;\kappa_{\mathrm{FE}}$ where
$\sup_{\begin{subarray}{c}w\in C\\ \|w\|=1\end{subarray}}w^{\prime}T(\beta)w\;\leq\;1-\eta_{C},\qquad\|P_{0}\,T(\beta)\|_{2}\;\leq\;\kappa_{\mathrm{FE}}.$
Assumption 8 says that usable variation on the positive-$Y$ part of the data dominates any fixed-effects-induced concentration: specifically, $\eta_{C}>\kappa_{\mathrm{FE}}$. The first bound ensures that, along any direction spanned by the doubly residualized regressors, the iteration weights never place all mass on boundary observations with $Y_{i}=0$; there is a uniform slack $\eta_{C}$ away from that boundary. The second bound limits how much these boundary weights can concentrate within fixed-effects groups; $\kappa_{\mathrm{FE}}$ captures the worst such concentration toward $Y_{i}=0$ within any group. Consequently, when some FE groups contain many zeros, the share of variation contributed by units with $Y_{i}>0$ must be correspondingly larger. In particular, groups composed entirely of zeros are ruled out.
The asymptotic properties are established in Theorem 4.
Theorem 4 (Consistency and Normality of iOLS-FD-HDFE)
Under Assumptions 1 2, and 8, the iOLS-FD-HDFE estimator is consistent and achieves the parametric rate of convergence $n^{-1/2}$. Formally, we have $n^{1/2}|\hat{\beta}_{t(n)}-\beta|=O_{p}(1)$ as $n\to\infty$ for any $t(n)\geq-\frac{1}{2}\log(n)/\log(\kappa)$, where $\kappa\in[0,1)$ is the modulus of the associated contraction mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$. In addition, iOLS-FD-HDFE is asymptotically normally distributed such that $\sqrt{n}\left(\hat{\beta}_{t(n)}-\beta\right)\overset{d}{\to}\mathcal{N}(0,\Omega)$, as $n\to\infty$, where the covariance can be consistently estimated using the asymptotic covariance of the OLS estimator in the last iteration up to minor modifications.
Therefore, the properties of iOLS established in Theorem 1 hence carry over to iOLS-FD-HDFE, provided FE groups in $X_{0}$ do not contain too many zero observations.
@@ 6 Conclusion
Appendix A Proof of Theorem 1: iOLS
Proving Theorem 1 requires multiple steps. We first prove consistency and normality for the biased case, i.e. $\forall\delta<\infty$. Then we prove consistency and normality for the exact case, i.e. $\delta=\infty$. Theorem 1 is just a sequential application of these results.
@@ Appendix A Proof of Theorem 1: iOLS
where $\tilde{Y}_{i}(\beta)=\log(Y_{i}+\delta\exp(X_{i}^{\prime}\beta))-c(\beta,\delta)$ is the transformed dependent variable. To further simplify exposition, we focus on $\delta=1$ in the proofs. The mapping from $\mathbb{R}^{K}$ to $\mathbb{R}^{K}$ which characterizes the parameter is hence defined $\forall\phi\in\mathbb{R}^{K}$ as
@@
Our proof follows Dominitz and Sherman (2005), hereafter denoted DS, who develop a convergence theory for iterative estimators. Following DS, the convergence of iOLS requires that $M(\cdot)$ and $\hat{M}_{n}(\cdot)$ be contraction mappings, asymptotically. [^27] The reader is referred to DS for a formal definition of an asymptotic contraction mapping. In order to show the convergence result $n^{1/2}|\hat{\beta}_{t(n)}-\beta|=O_{p}(1)$ as $n\to\infty$ by applying Theorem 1 in DS, we need to show that the following conditions hold:
@@
We verify conditions (i)-(iv) in turn. For (i), we adapt Dominitz and Sherman (2005) to show that $\{\hat{M}_{n}(\cdot)\}$ is an asymptotic contraction on $(B_{0},E_{K})$. In addition, we also show that, for any finite $n>0$ and any starting value $\phi$, $\hat{M}_{n}$ is a (finite-sample) global contraction. We report both results because Phase 2 of our algorithm (the final debiasing step) enjoys only a local contraction property, for which the asymptotic notion in Dominitz and Sherman (2005) is needed.
@@
Let us denote $X_{1i}=1$ for all $i$ as the constant. By prewhitening, we have $\sum_{j=1}^{n}X_{1j}=n$ and $\sum_{j=1}^{n}X_{kj}=0$ for $k>1$.
$\frac{\partial\hat{c}(\phi)}{\partial\phi_{k}}$ $=n^{-1}\sum_{i=1}^{n}\frac{\exp(X_{i}^{r^{\prime}}\phi^{r}+\hat{\phi}^{1})\,\big(\frac{\partial\hat{\phi}^{1}}{\partial\phi_{k}}+X_{ki}\big)}{Y_{i}+\exp(X_{i}^{r^{\prime}}\phi^{r}+\hat{\phi}^{1})}\;-\;n^{-1}\sum_{i=1}^{n}\big(\tfrac{\partial\hat{\phi}^{1}}{\partial\phi_{k}}+X_{ki}\big),$ (34)
for $k>1$, and $\frac{\partial\hat{c}(\phi)}{\partial\phi_{1}}=0$. This expression simplifies when evaluated at $\phi=\beta$:
@@
for $k>1$, because $\hat{\phi}^{1}(\beta)=\log\!\big(n^{-1}\sum_{i=1}^{n}Y_{i}\exp(-X_{i}^{r^{\prime}}\beta^{r})\big)=\beta_{1}+\log\!\big(n^{-1}\sum_{i=1}^{n}U_{i}\big)$ with $\log\!\big(n^{-1}\sum U_{i}\big)=O_{p}(1)$ by i.i.d. and $E[U_{i}]=1$, and $n^{-1}\sum_{i=1}^{n}X_{ki}=0$ by prewhitening. Thus, $\frac{\partial\hat{\phi}_{1}(\beta)}{\partial\phi_{k}}=0$.
Therefore, each element $(k,l)$ of $\hat{V}_{n}(\beta)$ is
$\left[\hat{V}_{n}(\beta)\right]_{k,l}$ $=n^{-1}\sum_{i=1}^{n}\frac{X_{ki}X_{li}}{1+U_{i}}\;-\;\Big(n^{-1}\sum_{i=1}^{n}X_{ki}\Big)\Big(n^{-1}\sum_{j=1}^{n}\frac{X_{lj}}{1+U_{j}}\Big),$ (36)
for $l>1$, and
$\left[\hat{V}_{n}(\beta)\right]_{k,1}$ $=n^{-1}\sum_{i=1}^{n}\frac{X_{ki}}{1+U_{i}},$ (37)
for $l=1$. Remark that for $k=1$ and any $l>1$ we have $\left[\hat{V}_{n}(\beta)\right]_{1,l}=0$, and for $k=1,l=1$, we have $\left[\hat{V}_{n}(\beta)\right]_{1,1}=n^{-1}\sum_{i=1}^{n}\frac{1}{1+U_{i}}<1$. Therefore, the eigenvalue associated with the constant term is strictly below $1$, and proving convergence amounts to showing that the largest eigenvalue of the $(K-1)\times(K-1)$ lower-right submatrix of $\hat{V}_{n}(\beta)$ is strictly less than unity. All elements $(k,l)$ for $k,l>1$ of this matrix are
$\left[\hat{V}_{n}(\beta)\right]_{k,l}$ $=n^{-1}\sum_{i=1}^{n}\frac{X_{ki}X_{li}}{1+U_{i}},$ (38)
@@
where $W=\mathrm{diag}\!\big(\tfrac{1}{1+U_{i}}\big)$ has diagonal entries in $(0,1]$. For general $\delta>0$ these weights are $\frac{\delta}{\delta+U_{i}}\in[0,1)$; hence the contraction modulus approaches $1$ as $\delta\to\infty$, so smaller $\delta$ yields faster numerical convergence.
Write $W=W^{1/2}W^{1/2}$ and note the equivalent representation $W=I_{n}-D$ with $D=\mathrm{diag}\!\big(\tfrac{U_{i}}{1+U_{i}}\big)\in[0,1)$. Then
$\left[\hat{V}_{n}(\beta)\right]_{k,l>1}$ $=n^{-1}X^{\prime}(I_{n}-D)X\;=\;I_{K-1}\;-\;n^{-1}X^{\prime}D^{1/2}D^{1/2}X.$ (40)
It follows that the maximum eigenvalue is equal to
$\lambda_{\max}\!\Big(\left[\hat{V}_{n}(\beta)\right]_{k,l>1}\Big)\;=\;\max_{\|a\|=1}\Big\{1-a^{\prime}X^{\prime}DXa\,\Big\}.$ (41)
Define $D_{Y}:=\mathrm{diag}(\mathbf{1}\{Y_{i}>0\})$ and $\eta_{*}(\beta,\delta):=\inf_{\,i:\,Y_{i}>0}\frac{U_{i}}{\delta+U_{i}}\in(0,1).$ Since $D\succeq\eta_{*}(\beta,\delta)\,D_{Y}$, we have
$a^{\prime}X^{\prime}DXa\;\geq\;\eta_{*}(\beta,\delta)\,a^{\prime}X^{\prime}D_{Y}Xa.$
Under Assumption 3 (overlap on positive outcomes), $a^{\prime}X^{\prime}D_{Y}Xa\geq\alpha\,a^{\prime}X^{\prime}Xa=\alpha$ by prewhitening. Therefore
$\lambda_{\max}\!\Big(\left[\hat{V}_{n}(\beta)\right]_{k,l>1}\Big)\;\leq\;1-\alpha\,\eta_{*}(\beta,\delta)\;<\;1.$
Thus, as $n\to\infty$, the maximum eigenvalue of $\hat{V}_{n}(\beta)$ is strictly less than unity, and the mapping is a contraction.
@@
where $W$ denotes the diagonal matrix with elements $\frac{\exp(X_{i}^{\prime}\phi)}{Y_{i}+\exp(X_{i}^{\prime}\phi)}\in(0,1]$. Therefore, the same logic applies and the algorithm is a contraction mapping for any $n$ and any $\phi$, under the overlap condition.
@@
Following the same reasoning as in the previous theorem, a sufficient condition for convergence is that $\frac{U_{i}}{1+\rho}$ is between 0 and $1$ for all $i$, as well as an overlap on positive outcomes condition. Therefore, the choice of $\rho$ will not only affect the speed of convergence, but also whether the estimator converges at all.
@@
has element $1-\frac{U_{i}}{1+\rho}$, and $\hat{\Sigma}_{0}$ is an estimator of the covariance of $X_{i}^{\prime}U_{i}$ across observations.
Appendix B Proof of Theorem 2: i2SLS